If you have 150grams of s8, how many more of so2 can you make?
1
Expert's answer
2019-12-11T07:54:26-0500

S8 = 8SO2

n(S8) = m/M = 150g/256g/mol = 0.586 mol

n(SO2) = 8*n(S8) = 8*0.586 = 4.688 mol

m(SO2) = n*M = 4.688mol*64g/mol = 300 g